OpenStudy (anonymous):

Lasers are now used in eye surgery. Given the wavelength of a certain laser is 514 nm and the power of the laser is 1.6 W, how many photons are released if the laser is used for 0.058 s during the surgery?

OpenStudy (ash2326):

Energy of a photon =hv h=6.626*10^-34 v=frequency=3*10^8/*(514*10^-9) E =3.8673*10^-17 Joules now lase operates for 0.058 seconds Power=1.6w Energy released in this time=1.6*0.058 =0.0928 Joules No. of Photons=Total Energy released/energy of a photon =0.0928/(3.8673*10^-17) =0.02399*10^17 no. of photons released =2.399*10^15

OpenStudy (anonymous):

Off by a power of ten but none the less the reasoning is right!